### v1

# Class-Discriminative Attention Maps for Vision Transformers

Abstract
Interpretability methods are critical components for examining and exploring deep neural networks (DNN), as well as increasing our understanding of and trust in them. Vision transformers (ViT), which can be trained to state-of-the-art performance with a self-supervised learning (SSL) training method [1], provide built-in attention maps (AM). While AMs can provide high-quality semantic segmentation of input images, they do not account for any signal coming from a downstream classifier. We introduce class-discriminative attention maps (CDAM), a novel post-hoc explanation method that is highly sensitive to the target class. Our method essentially scales attention scores by how relevant the corresponding tokens are for the predictions of a classifier head. Alternative to classifier outputs, CDAM can also explain a user-defined concept by targeting similarity measures in the latent space of the ViT. This allows for explanations of arbitrary concepts, defined by the user through a few sample images. We investigate the operating characteristics of CDAM in comparison with relevance propagation (RP) [2] and token ablation maps (TAM), an alternative to pixel occlusion methods. CDAM is highly class-discriminative and semantically relevant, while providing implicit regularization of relevance scores.
PyTorch implementation: https://github.com/lenbrocki/CDAM.
Web live demo: https://cdam.informatism.com/.
Figure 1: Extending attention maps (AM), the proposed class-discriminative attention maps (CDAM) quantify and visualize input features that are relevant for the target class in Vision Transformers (ViT). We trained a linear classifier on the ViT-S/8 trained with DINO as a backbone [1] and visualize random samples with CDAM for different target classes. For details, refer to Section 3.1. Orange and blue colors correspond to positive and negative values, respectively.
1 Introduction
The transformer architecture, which was originally developed for natural language processing [3], has been developed into vision transformers (ViT) that demonstrate comparable or even better performance than convolutional neural networks (CNN) on many computer vision tasks [4, 5]. Additionally, a powerful feature of transformers is that the attention mechanism provides a built-in explanation of the model’s inner workings, if attention is understood as a weighing of the relevance of individual tokens [6, 7, 8, 9]. While attention maps (AM) provides intuitive interpretability, it fails to be discriminative with respect to a target class. Recognizing the potential of AM for explaining a ViT’s inner workings, we propose class-discriminative attention maps (CDAM), a gradient-based extension that makes them sensitive to class-specific signals coming from a downstream classifier.
Recently, it has been demonstrated that ViT models trained with DINO, a self-supervised learning (SSL) training method [1], produce AMs that provide high-quality semantic segmentations of input images. These AMs clearly distinguish objects of interest from the background with very little noise, with different attention heads often focusing on semantically distinct regions [1]. AMs are the visualization of the self-attention of the [CLS] token query in the final attention layer. However, one shortcoming of AMs is that they purely operate on the level of the last attention layer and therefore do not take into account any information coming from a downstream task, such as a classifier head.
The proposed CDAMs rely on the calculation of gradients of the class score, i.e. the logit value of a target class, with respect to the token activations in the final transformer block (Fig. 2(a)). We show through explicit calculation and empirical observations that CDAM scales the attention by how relevant the corresponding tokens are for the model’s decision. CDAM therefore inherits the desirable properties of AMs, but additionally indicates the (counter-)evidence for a target class that the ViT relies on for its predictions (Fig. 1). We demonstrate that it is class-discriminative in the sense that the heat maps clearly distinguish the object representing the target class from the background and, importantly, from objects that belong to other classes. CDAMs are also highly sensitive to the target class which, all together, makes it a useful tool for explaining the decision-making process of classifiers with a ViT backbone.
In addition to the classifier-based explanations, we present a concept-based approach that replaces the class score with a similarity measure in the latent space of the ViT, where a concept is defined through a number of example images. Averaging the latent representations of the example images yields a concept embedding, which is analogous to a concept vector in variational autoencoders (VAE) and related generative models [10, 11]. In brief, we use CDAM to identify image patches that drive the (dis-)similarity of the concept embedding and the latent representation of a given image. This allows obtaining a heat map for the target concept the model has never been explicitly trained on, which is of particular interest for self-supervised models. Both supervised and semi-supervised variants of CDAM are compared to token ablation maps (TAMs) and relevance propagation (RP) [2]. Overall, our results suggest that CDAMs exhibit semantic coherency, implicit regularization, and strong discrimination with respect to a target class/concept.
2 Related Works
Interpretability of ViT, and more broadly deep neural networks (DNN), is an active area of research, since understanding their decision-making process would help improve the model, identify weaknesses and biases, and ultimately increase human trust in them. The proposed class-discriminative attention map (CDAM) is related to both attention and gradient-based explanation methods for deep learning models. On the one hand, explanations are obtained via backpropagation of gradients, but on the other hand, we show that the CDAM scores are directly proportional to the attention.
Gradient-based methods operate by backpropagating gradients from the prediction score to the input features to produce feature attribution maps, also called saliency maps. Many variants have been proposed, see [12, 13, 14, 15, 16] to just name a few. In particular, CDAM is similar to and inspired by the input×$\times$×gradient method [16], because it involves calculating the gradients with respect to tokens multiplied by their activation (Eqs. 2 and 4). The main difference to the input×$\times$×gradient method, and most of other gradient methods, is that we don’t backpropagate the gradients all the way to the input tokens. CDAM stops backpropagation at the tokens that enter the final attention layer. In that regard, our method is related to GradCam [14] which backpropagates the gradients to the final convolutional feature map in a CNN. GradCam and CDAM therefore share the property that they operate on high-level features which is presumably the reason why they are both sensitive to the targeted class, a feature that many explanation methods do not have [17]. Several methods have attempted to apply gradient-based methods or Layer-Wise Relevance Propagation (LRP) [18] to Transformer architectures [19, 20, 21, 22, 23], mostly in the context of NLP tasks.
Attention has often been used to gain insight into a model’s inner workings [6, 7, 8, 9] since it seems intuitive that this weighing of individual tokens correlates with how important they are for the performed task. There is also an ongoing discussion about whether attention provides meaningful explanations [8, 24, 25]. Furthermore, methods such as “attention rollout” [22] have been proposed that aim to quantify how the attention flows through the model to the input tokens. [2] defines the relevancy matrices, as the Hadamard product of the attention map and the gradient of $f_{c}$ w.r.t. to the attention layer, accumulated at each layer. This relevance propagation (RP) method [2], to the best of our knowledge, shows the highest degree of class-discrimination in ViT models. This is used in our study as a comparison for our method.
Many papers have been published on the problem of evaluating post-hoc explanation methods, leading even to the proposal of meta-evaluations [26]. In a recent overview of quantitative evaluation methods [27] 12 desirable properties have been identified. We primarily focus on three of them, namely correctness, contrastivity, and compactness. Our evaluation of correctness is a variant of Sensitivity-$n$ [28], which repeatedly masks $n$ randomly selected pixels and compares the sum of their relevance scores with the change in model output. In the context of ViTs, instead of arbitrarily perturbing observed pixels to some non-informative values, our method called token ablation map (TAM) removes patch tokens before they enter the first transformer block, followed by measuring the resulting difference in model performance. We consider TAM as an intuitive first-order approximation that serves as a sanity check for CDAM. Compactness is measured by looking at the shrinkage and sparsity of CDAM compared to other methods.
3 Proposed Methods
Figure 2: Graphical scheme of class-discriminative attention maps (CDAM). (a) CDAMs are obtained by first propagating an image through the transformer blocks and using its latent representation $l=\texttt{[CLS]}^{\prime}$ to infer a class score $f_{c}$ or similarity score $g(l,l_{c})$. The gradients wrt the layer-normalized input tokens $T_{i}$ of the final attention layer and their activations are then used to calculate the relevance score $S_{i,c}$ according to Eq. 2 or Eq. 4. (b) Detailed view of the final transformer block, adapted from [4].
[…]
In the first scenario, the class score is obtained from the corresponding activation in the prediction vector of a classifier trained on top of the ViT in a supervised fashion. In the second scenario for the concept similarity score, we define a concept through a small set of examples (e.g. ten images of a dog) and obtain a concept vector by averaging their latent representations. Then, without training a classifier, a similarity measure between the concept vector and the latent representation of a targeted sample image is computed and used as the target for the gradients. We call the resulting feature attribution maps “class-discriminative attention maps (CDAM)” because we observe that they scale the attention scores
$A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right),$ (1)
by how relevant the corresponding tokens $T_{i}$ are for the class or similarity score. $Q$ and $K$ are the query and key matrices, consisting of a set of $d_{k}$ dimensional vectors [3]. We discuss the proportionality of CDAM to these attention scores in Section 5.1.
3.1 Relevance Score from Classifiers
[…]
$S_{i,c}=\sum_{j}T_{ij}\frac{\partial f_{c}}{\partial T_{ij}},$ (2)
where $T_{ij}$ is the $j$-th component of the $i$-th token that is fed to the last attention layer [^1] We found in our experiments that calculating the gradients wrt the layer-normalized tokens yields far better results than using the unnormalized ones. (Fig. 2). This definition is a first-order estimation of the contribution of $T_{i}$ to $f_{c}$ in the sense that it would be exact if $f_{c}$ depended on $T_{ij}$ only linearly. It is inspired by gradient-based explanation methods with the difference that we don’t backpropagate the gradients all the way to the input layer but only to the input tokens to the final attention layer. This allows us to gain insight into the decision-making process of the combination of ViT and classifier based on high-level features.
This approach leads to attention maps that are class-discriminative and well-aligned with the semantic relationships of the class and different parts of an image (Fig. 1). As a side note, the heatmaps shown in this article are all scaled independently, thus there is no meaningful comparison of the scale of relevance between images.
3.2 Relevance Score from Similarity Measures
Another way of probing high-level features learned by the ViT is through the use of a similarity measure $g$ in the latent space of the ViT. We compare the latent representation $l(X)=T_{\texttt{[CLS]}^{\prime}}(X)$ of a sample image $X$ and a concept vector $l_{c}$ [29, 10, 11], i.e. the averaged latent representation of $n$ images that are chosen to represent the concept of interest:
$l_{c}=\frac{1}{n}\sum_{X\in c}l(X).$ (3)
This approach is directly inspired by the concept saliency map [10]. The definition of the relevance score in this setting is analogous to Eq. 2
$S_{i,c}=\sum_{j}T_{ij}\frac{\partial g(l,l_{c})}{\partial T_{ij}}.$ (4)
Possible choices for $g$ are for example the $L^{2}$ distance, cosine similarity or dot product, where we use the latter in our experiments. The definition Eq. 4 estimates the contribution of $T_{i}$ to $g$, i.e. it allows us to gauge how much each token is driving the (dis-)similarity between $l$ and $l_{c}$.
3.3 Token Ablation Study
As an independent comparison method, we carried out a token ablation study, in which a change in the model accuracy is measured after removing one token at a time. We visualize changes in the model accuracy as token ablation maps (TAM). This method for ViT is analogous to input pixel occlusion [30]. Such occlusion methods gauge a (group of) pixel’s relevance by masking them with non-informative values and recording the change in model output. This approach is problematic because perturbing pixels potentially leads to out-of-distribution samples which may trigger artifacts of the model [31, 32]. Relatedly, in contrast to pixel occlusion methods, for TAM there is no need to choose a specific way of perturbing input features in pixel space, which would lead to some ambiguity and variation in the resulting relevance maps. The token ablation study sidesteps this issue in the context of ViTs by removing image patch tokens, one at a time, before they enter the first transformer block, instead of masking image patches in pixel space. The relevance score assigned by TAM is defined by
$S^{\text{TAM}}_{i,c}=f_{c}(\mathbf{T})-f_{c}(\mathbf{T}\backslash{T_{i}}),$ (5)
where $\mathbf{T}$ is the set of all tokens that enter the first transformer block and $\mathbf{T}\backslash{T_{i}}$ has token $T_{i}$ removed from $\mathbf{T}$.
Note that TAMs are not to be confused with the “ground truth” of feature relevance since we are dealing with non-linear models with complex interactions of patch tokens. Therefore, tokens would likely be assigned different relevance scores if we perturbed two or more tokens simultaneously, with different scores depending on the chosen group of tokens. This effect has previously been demonstrated in the context of CNNs [28]. Therefore, comparing with TAM mainly serves as a sanity check for CDAM, since TAM is a first approximation to patch token relevance with a clearly defined meaning.
4 Results
Figure 3: Class-Discriminative Attention Maps (CDAM) for user-defined concepts (Section 3.2). The concept embedding $l_{c}$ has been obtained by averaging the latent representations of 30 images that include the common concept. In each instance, from left to right: Sample image, AM, and CDAM. Orange and blue correspond to positive and negative values, respectively.
For our experiments, we have been using the ViT-S/8 architecture trained with DINO as a backbone [1], which is available as a pre-trained model in their repository [^2] https://github.com/facebookresearch/dino.
4.1 Relevance Scores from Classifier
To demonstrate our method, we have trained a linear classifier on the ImageNet [33] that achieved an accuracy of $77.0\%$. We use random resized cropping and horizontal flipping with PyTorch default arguments as augmentation, Adam optimizer with learning rate $3\times 10^{-4}$, batch size 128, and train for 10 epochs. The parameters of the ViT backbone are frozen during training, so only the classifier head is trainable. The predictions of this classifier were then used to create the heat maps shown in Fig. 1 via Eq. 2. Notice how they clearly distinguish the targeted class with positive values (evidence) from the rest of the image, which is assigned either zero (irrelevant) or negative scores (counter-evidence). Note that CDAM does not assign non-zero $S_{i,c}$ to tokens that have zero attention in AM. CDAM therefore appears to compute $S_{i,c}$ by multiplying the attention with the relevance of the corresponding token for the targeted class. This is a nice feature since CDAM inherits the high-quality object segmentations present in AM, which are focused on the objects of interest. We discuss this operating characteristic in Section 5.
One of the most interesting features of CDAM is that it is highly sensitive to the chosen class (Fig. 4, left). For example, when either the zuccini or bell pepper class is chosen, CDAM clearly highlights that vegetable. Most of the other vegetables are assigned either near-zero or negative values. Interestingly, the green stems of other vegetables are also assigned positive values. CDAM reveals that the ViT has learned to associate input features with chosen classes that are largely consistent with human expectations. Additional examples of CDAM and other interpretability methods are showned in the Appendix A (Figs. 7, 8, 9 and 10).
Figure 4: When targeting different classes (left) or concepts (right), the resulting CDAMs are distinct and align with the corresponding objects. Note that green parts of vegetables are shown to activate the zuccini class (left). The wrinkles in the elephant’s trunk appear to be related to the zebra concept, probably semantically mistaken by the model as zebra stripes (zoom in for a better view).
4.2 Relevance Scores from Similarity Measure
Concept embeddings $l_{c}$ have been obtained by averaging 30 latent representations of images depicting the concept of interest. These images were randomly selected from the corresponding classes in the ImageNet validation set [33] and the CDAM in Fig. 3 were obtained using Eq. 4. Again we observe that CDAM reveals that the model clearly distinguishes the parts of the images that are relevant for the chosen concept from the rest of the image. For instance, for the concept hammer, CDAM highlights not only the hammer itself but also the wooden handles of the screwdrivers and the top of the ratchet, which are indeed relevant for this concept. Unrelated tools or parts appear in blue indicating that the model deems them to be not similar to the concept hammer.
Just as CDAM is sensitive to the chosen class, it also produces attention maps that are specific to the chosen concept (Fig. 4, right). This makes the proposed method particulary valuable for understanding the representations the ViT has learned since the explanations for different concepts can give a complementary view. For instance, CDAM shows that parts of the elephant’s trunk are counter-evidence for the concept african elephant (Fig. 4, right) which is at first counter-intuitive. However, when compared with the concept zebra, we can see that the model deems this part of the trunk to be similar to a zebra and it probably mistakes the trunk’s wrinkles for the zebra’s stripes. By choosing a set of images with a common concept, CDAM helps probing the latent space of the ViT, allowing us to learn about how it “sees” the world in a semi-supervised manner.
4.3 Comparison with relevance propagation and token ablation maps
Figure 5: Comparison of specificity to the target class. AM is, by definition, not influenced by it, but relevance propagation [2] does show some specificity. In particular due to the signed relevance scores both CDAM and TAM exhibit strong class-sensitivity and the two maps roughly agree, with CDAM appearing as stronger regularized.
Upon visual inspection, it appears that CDAM is estimating patch token relevance that aligns well with our expectation concerning the most relevant (counter-)evidence for a given target class (Figs. 1 and 4). To investigate whether these relevance scores correspond to the actual inference performed by the model, we compare them to token ablation maps (TAMs) and relevance propagation (RP) [2]. In TAMs, the model output is measured after an individual token is removed (Section 3.3). RP, which is based on the Deep Taylor Decomposition, is an explanation method for ViTs that, to the best of our knowledge, produces relevance scores that are the most class-discriminative so far [2, 23].
We find that RP is more class-discriminative than AM and that CDAM and TAM distinguish the different classes more clearly than RP ( Fig. 5). Visually, the signed scores in CDAM and TAM make stronger distinctions between the different classes, whereas RP only “dims” the object that is not targeted. Fig. 6 shows pair-wise comparisons among scores for CDAM, TAM, AM, and RP. We observe that RP tends to assign low but positive scores to tokens that obtained negative scores from CDAM or TAM.
[…]
Compared to AM, TAM, and RP, CDAM appears to be less noisy (Fig. 5). We can observe this implicit regularization effect by looking at the scores from 1000 randomly selected images Fig. 6. In AM, TAM, and RP, the per-image average percentage of patch tokens with relevance scores smaller than $5\%$ of the maximum score are $54\%,66\%,49\%$, respectively, whereas in CDAM it is $84\%$. While sparsity may not always imply the most accurate estimation at a single data point (e.g., a pixel), regularization has long been pursued and motivated by interpretability in machine learning [34].
Figure 6: Pairwise scatter plots of relevance/attention scores for CDAM, TAM, AM, and relevance propagation obtained from 1000 randomly selected images from the ImageNet validation set. The target class for the explanations is the one with the highest logit score. Notice in particular that AM and relevance propagation are highly correlated and that CDAM and TAM largely agree on the sign of relevance scores. We also observe that $A_{i}=0\Rightarrow S_{i,c}=0$, where $A_{i}$ and $S_{i,c}$ refer to the AM or CDAM score of token $T_{i}$.
5 Discussion
Class-Discriminative Attention Map (CDAM) is a gradient-based extension of attention maps (AM), that provides strong discrimination with respect to a chosen class or concept. In contrast, many post-hoc explanation methods produce similar heatmaps even when targeting different classes [23, 17]. CDAM retains appealing characteristics of AM, namely high-quality semantic segmentation in self-supervised ViTs [1]. Previously, a considerable effort has been made to de-noise post-hoc interpretability methods by averaging variations of multiple saliency maps [13, 15]. Without further processing or averaging, the proposed CDAM produces comparatively sparse heat maps that are less noisy than previous methods.
Besides explaining predictions of the classifier on the top of ViT, we also introduce how to obtain a CDAM by measuring the similarity between the latent representation of a sample image and a concept embedding. By probing the latent space of the ViT with different concepts (Fig. 4) we can learn a lot about the representations it has learned. To some degree this can of course also be done by looking at the attention maps of the different heads, which tend to specialize in different categories of objects or parts thereof [1]. This approach is, however, limited to the categories that the attention heads happened to have learned. CDAM allows us to probe for concepts that are relevant to any given use case.
5.1 Proportionality of CDAM to the attention
CDAM appears to scale the attention $A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right)$ by the relevance of the corresponding token $T_{i}$ for a concept. To better understand this observation, without loss of generality, consider a single attention head. Eq. 4 is expanded
$S_{i,c}$ $=\sum_{j}T_{ij}\frac{\partial g(l,l_{c})}{\partial T_{ij}}$
$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}l_{k}l_{c,k}$
$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}h_{k}\left(\sum_{m}A_{m}V_{m}+\texttt{[CLS]}\right)l_{c,k}$
$=\sum_{j}T_{ij}\sum_{k,m,n}\nabla_{n}h_{k}\times\left(\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}+A_{m}\frac{\partial V_{mn}}{\partial T_{ij}}\right)l_{c,k}$
$=\sum_{j,k,m,n}T_{ij}\nabla_{n}h_{k}\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}l_{c,k}+\sum_{k,n}A_{i}V_{in}\nabla_{n}h_{k}l_{c,k},$ (6)
[…]
Importantly, we observe in our multi-head attention ViT model (Fig. 6) that
$A_{i}=0\Rightarrow S_{i,c}=0$ (7)
which implies that in Section 5.1, all terms that are not proportional to $A_{i}$ are zero or cancelled out and we are left with
$S_{i,c}=A_{i}V_{i}\cdot\nabla(h\cdot l_{c}).$ (8)
Empirically, $S_{i,c}$ is obtained by multiplying $A_{i}$ with the directional derivative of $h\cdot l_{c}$ in the direction of $V_{i}$. In other words, CDAM scales $A_{i}$ by the rate at which the dot product (similarity) between $h$ and $c$ changes in the direction of $V_{i}$. In the case that $h$ is just the identity function, we obtain an even simpler interpretation of CDAM. That’s because we then have $\nabla_{n}h_{k}(f)=\partial_{f_{n}}f_{k}=\delta_{nk}$ and get from Section 5.1, again assuming $\frac{\partial A_{m}}{\partial T_{ij}}=0$, that $S_{i,c}=A_{i}V_{i}\cdot l_{c}$, which means that $A_{i}$ is scaled by the similarity of $V_{i}$ and $l_{c}$.
5.2 Faithfulness of Post-hoc Explanation Methods
Understanding the decision-making process of deep neural networks is intrinsically challenging due to the vast parameter space and non-linearity, which lead to complex interactions among the patch tokens in ViT. Therefore, post-hoc explanation methods such as RP, CDAM, and TAM can only explain some facets of the model and can never be comprehensively faithful to it, as that would entail being equivalent to the model itself [17]. Of course, the complete set of the model weights in DNNs would not be considered interpretable. Therefore, attempting to create explanation methods necessitates a balance among many properties, ranging from sparsity to discriminativeness [27].
The proposed CDAM relies on gradients and is therefore a first-order approximation. Instead of referring directly to the input patch tokens, the CDAM scores refer to tokens that, in general, have been enriched with contributions from all other tokens through the attention layers. Relevance is therefore not directly attributed to input features. Nonetheless, by extending the inherent attention scores in ViTs, CDAM captures the high-quality semantic segmentations of input images shown in self-supervised ViTs [1], while providing class-discrimination.
6 Conclusion
By extending the attention maps (AM) inherent to ViT, the proposed CDAM provides relevance scores for a chosen class or concept. The proposed relevance estimation for CDAM essentially scales the attention scores by how relevant they are for a target class or concept. As a result, zero attention scores remain zero in CDAM. The resulting heat maps clearly distinguish the object representing the target from the background, as well as other objects that are present. In addition to using a classifier, we proposed to explain a user-defined concept by selecting a small set of example images. Using similarity measures in the latent representation, CDAM provides highly specific heat maps for a concept that underly selected examples. Especially in the context of SSL where class labels are absent, this makes it a valuable technique to investigate the model representations.
In terms of class-discrimination, our method compares favorably against the state of the art, namely relevance propagation (RP) for transformers [2]. Especially, the positive and negative signs of relevance scores in CDAM make clear distinctions qualitatively and quantitatively. We also devise the token ablation maps (TAM) which visualize the changes in the ViT model output when corresponding image patch tokens are removed. The advantage of TAM is its explicable computation, which serves as a sanity check. CDAM captures the overall patterns of TAM while providing sparser and more semantically consistent heat maps. However, while RP and TAM are used in comparison, neither of them is the “ground truth” of relevance scores.
Post-hoc explanation methods for DNNs approximate the full operating characteristics of the model with certain desirable properties. Our results suggest that CDAM is capable of achieving a high degree of correctness, contrastivity, and compactness, as described in [35]. Note that CDAM is not explicitly designed to satisfy these criteria. For example, we could impose further sparsity constraints that would increase compactness. However, over-optimizing an explanation method for certain properties could have undesirable effects [^3] Goodhart’s Law states that “when a measure becomes a target, it ceases to be a good measure” [36]. While this paper uses a ViT model trained with SSL [1], CDAM is applicable more broadly to all transformer models with self-attention, e.g., large language models [37].
Acknowledgments
This work was funded by the ERA-Net CHIST-ERA grant [CHIST-ERA-19-XAI-007] long term challenges in ICT project INFORM (ID: 93603), by the National Science Centre (NCN) of Poland [2020/02/Y/ST6/00071]. This research was carried out with the support of the Interdisciplinary Centre for Mathematical and Computational Modelling University of Warsaw (ICM UW) under computational allocation no GDM-3540; the NVIDIA Corporation’s GPU grant; and the Google Cloud Research Innovators program.
Appendix A Appendix
A.1 Multiple attention heads
We note that introducing $h$ attention heads does not change the derivations in Section 5.1 qualitatively. Instead of the sum $Z=\sum_{m}A_{m}V_{m}$ there will be a term $\text{Concat}(Z^{1},Z^{2},..,Z^{h})W^{O},W^{O}\in\mathbb{R}^{hd_{v}\times d_{\text{model}}}$ so that in the case of multi-head attention the equivalent of Section 5.1 reads
$S_{i,c}=\text{Concat}\left(Y^{1},Y^{2},..,Y^{h}\right)W^{O}\nabla(h\cdot l_{c}),\;Y^{n}=T_{i}\sum_{m}\frac{\partial A^{n}_{m}}{\partial T_{i}}V^{n}_{m}+A^{n}_{i}V^{n}_{i}$ (9)
and since $A^{n}_{i}>0,A_{i}=\sum_{n}\frac{A_{i}^{n}}{h}$ implies that $A^{i}=0\Rightarrow A_{i}^{n}=0,\forall n$ we can again conclude that in order for Eq. 7 to hold all terms that are not proportional to $A_{i}$ either vanish or cancel out. For multi-head attention we thus arrive at
$S_{i,c}=\text{Concat}\left(Y^{\prime 1},Y^{\prime 2},..,Y^{\prime h}\right)W^{O}\nabla(h\cdot l_{c}),\;Y^{\prime n}=A^{n}_{i}V^{n}_{i},$ (10)
which does not have such a straightforward interpretation as the single-headed case due to the mixing of the attention heads.
If we think of $W^{O}$ as acting first on $\nabla(h\cdot l_{c})$, creating a column vector of dimension $hd_{v}$, we can understand Eq. 10 as the sum of $h$ directional derivatives of the $d_{v}$-long segments of this columns vector in the directions of $Y^{\prime 1}$ to $Y^{\prime h}$. Each of the summands is proportional to the corresponding $A^{n}$, and the interpretation of CDAM from the single-head case therefore extends to the multi-head one.
Although our observation $A_{i}=0\Rightarrow S_{i,c}=0$ implies that all terms that are not proportional to $A_{i}$ in Section 5.1 should vanish, we don’t have a theoretical argument why that would be the case. It seems unlikely that the various terms in the first sum cancel out, which leaves $\frac{\partial A_{m}}{\partial T_{ij}}=0$ as an explanation. However, in general, $A_{i}$ depends on $T_{i}$ through $K_{i}$ and one would therefore expect the derivative to be non-zero.
We can also show that CDAM scales the attention by the relevance of the correpsonding token for a chosen class (instead of concept), in the case that a classifier with weights $W^{C}$ has been trained on top of the ViT. Analogously to the above single-head discussion we have
$S_{i,c}$ $=\sum_{j}T_{ij}\frac{\partial f_{c}}{\partial T_{ij}}$
$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}l_{k}W^{C}_{kc}$
$=\sum_{j,k,m,n}T_{ij}\nabla_{n}h_{k}\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}W^{C}_{kc}+\sum_{k,n}A_{i}V_{in}\nabla_{n}h_{k}W^{C}_{kc},$ (11)
and the only difference is that the directional derivative now acts on $\sum_{k}h_{k}W_{kc}$, i.e. the contribution of $h$ to the activation of class $c$ in the prediction vector, instead of $\sum_{k}h_{k}l_{c,k}$. All the intuitions we have gained from the concept-based CDAM therefore extend to the class-based one.
A.2 Examples for explanation methods
Figure 7: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure 8: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure 9: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure 10: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.